5. Order: D NS IV with 20 mEq KCI per liter to run at 25 ml/h Drop factor: 60 gtt/mL gtt/min
Answer #1

Ans)5) Formula:

ml/hr = Total volume to be Infused × drop factor/ time in hr

25 = 500 × 60/ x

x= 30000/25 = 1200

1200/60 = 20 gtt/min